OpenStudy (turingtest):

Ahoy, matey! A raft and a motorboat left town A simultaneously and traveled downstream to town B. (The raft always moves at the same speed as the current, which is constant.) The motorboat arrived at town B, immediately turned back, and encountered the raft two hours after they had set out from A. How much time did it take the motorboat to go from A to B? (Assume that it travels at a constant rate of speed.)

OpenStudy (anonymous):

they sail on river?

OpenStudy (turingtest):

well it's not sailing with a motorboat and a raft... no sail involved o-0 don't see how that makes a difference though.

OpenStudy (anonymous):

idk maybe word sailing is not appropriate but like car rides on road, so raft and motorboat sails on river? and if it's river we need to take account river speed or maybe it's lake so then we do not need

OpenStudy (turingtest):

well, all I'll say is that there is a slightly tricky math way to solve this, but if you can apply physics the problem is completely obvious...

OpenStudy (turingtest):

gotta go, I'll see what we have when I get back from class.

OpenStudy (anonymous):

so i don't understand where they are sailing? or it's unknown?

OpenStudy (turingtest):

to town "B", wherever that is... that doesn't affect the problem, we just want to know how much time it took the motorboat to get there.

OpenStudy (anonymous):

what does current means then?

OpenStudy (anonymous):

after googling it seems that you need to add current when sailing downstream and subtract viceversa so it means it's river and current is river speed, no?

OpenStudy (anonymous):

x - current speed y - motor boat speed t1 - time from A to B t2 - time from B to meeting point \[\begin{cases}t_1+t_2=2\\xt_1+xt_2+(y-x)t_2=(y+x)t_1\end{cases}\] now you solve this and get answer